# Dual-Model Weight Selection and Self-Knowledge Distillation for Medical Image Classification

### v1

@@ 1 Introduction
Deep learning has become indispensable in the medical field, driving significant advances in diagnostic imaging, disease prediction, and electronic health record analysis [1]. Large-scale models have achieved remarkable success across diverse tasks; however, their deployment in clinical settings is often hindered by limited computational resources [2, 3]. For example, analyzing high-resolution medical images such as chest X-rays and magnetic resonance imaging (MRI) scans requires substantial computing power and memory, making large model deployment impractical in many healthcare environments [4, 5]. Simultaneously, the growing shortage of healthcare professionals is increasing the need for efficient automated systems to reduce workloads [6]. Consequently, there is a rising demand for compact models that can operate effectively under resource constraints while maintaining performance comparable to large-scale counterparts.
A critical factor in improving small-model performance is weight initialization. Poor initialization can lead to vanishing or exploding gradients, hindering effective training. While commonly used methods such as Xavier [7] and Kaiming [8] initialization support stable training from scratch, pretrained large-scale models (e.g., ImageNet-21K [9] or LAION-5B [10]) are increasingly adopted for transfer learning. However, fine-tuning these large models is often infeasible in clinical settings because it demands high-end GPUs with large memory and substantial computing resources—conditions rarely met in real-world medical environments. To address this, recent studies have explored weight selection methods, which transfer only a subset of weights from a large pretrained model to a smaller one [11, 12].
@@ 2.2 Knowledge Distillation
SKD eliminates the need for an external teacher by enabling the model to improve itself. Zhang et al. [41] introduced a method for segmenting deep models such as ResNet into blocks, allowing intra-model knowledge transfer between components. Kim et al. [42] introduced progressive SKD, and Ge et al. [43] developed batch-wise ensembling of self-predictions to suppress incorrect predictions. Recent work has adapted these ideas for transformers [44] and even explored self-distillation in the absence of ground-truth labels using self-supervised techniques [45]. MiniLM [46] exemplifies a task-agnostic SKD method that maintains performance while significantly reducing the model size. A comprehensive survey on knowledge distillation summarizes these trends and outlines the future directions [47].
@@ 3 Dual-model Weight Selection and Self-knowledge Distillation
Our method employs a pretrained large model as the teacher and small-scale models as students, integrating dual-model weight selection for effective knowledge transfer and SKD for efficient training. First, weight selection transfers essential parameters from the teacher model to the student models, enabling them to inherit prior knowledge and achieve strong early performance. Then, SKD allows the student models to refine their learning by leveraging their own predictions, improving task performance, and maximizing the benefit of the initial weights without imposing excessive computational cost.
@@ 3.2 Self-knowledge Distillation
where $\eta$ denotes the learning rate, and $\theta_{S}$ represents the parameters of the main student model $S$. Then, the weights of the auxiliary student model $S^{\prime}$ are updated using an exponential moving average (EMA), in which the parameters of model $S$ are integrated into $S^{\prime}$, with the stop gradient (SG) operation applied. EMA and SG are applied during training to stabilize the learning process and reduce update variance. After each training iteration, the weights of the model $S^{\prime}$ are updated using:
@@ 4.1 Dataset and Settings
4.2 Main Results
Tables 2, 3, and 4 present the classification results for chest X-ray, lung CT scan, and brain MRI images, respectively. Each experiment was repeated five times, and both the mean and standard deviation were recorded. These tables compare the performance of CM1, CM2, CM3, and CM4 with the proposed method, which combines weight selection and SKD (Ours).
Our method consistently outperformed all comparison methods in classification accuracy across all datasets, as shown in Tables 2, 3, and 4. By integrating dual-model weight selection with SKD, it enables efficient knowledge transfer and reuse, significantly improving feature retention and learning efficiency. Compared to CM1, which applies weight selection alone, our approach significantly improves feature retention and learning efficiency. Moreover, by evaluating different initialization strategies, we observed that our method substantially enhances neural network initialization, yielding more effective training than CM2, CM3, and CM4. The consistent superiority across datasets demonstrates the robustness and generalizability of our approach, even in low-data regimes. Notably, the significant improvements over CM2–CM4 emphasize that default or classical initialization schemes alone are insufficient for effective learning in smaller models. Our method not only boosts accuracy but also ensures stability across trials, as reflected in the low standard deviations reported.
Figures 3, 4, and 5 illustrate the confusion matrices of our method on the chest X-ray, lung CT scan, and brain MRI datasets, respectively. The results demonstrate consistently high accuracy and robust generalization across all modalities. In the chest X-ray dataset (Fig. 3), the model accurately distinguishes COVID-19 from other classes, between clinically similar conditions such as viral pneumonia and lung opacity, underscoring its robustness even with limited data. In the lung CT dataset (Fig. 4), the model shows accurate classification, especially between COVID-19 and normal cases, which are often difficult to separate due to radiological differences. For the brain MRI dataset (Fig. 5), the model effectively identifies tumor types such as glioma and meningioma, while reducing misclassification involving the no tumor class. These results confirm that combining dual-model weight selection with SKD enhances feature representation and discrimination, even when trained on just 10% of the available data.
4.3 Analysis of Interchanged Main and Auxiliary Models
4.4 Analysis of Hyperparameters
@@
We conducted experiments using ViT-S as the teacher model and ViT-T as the student model to analyze the effects of three key hyperparameters: $\alpha$, $\beta$, and $\tau$. The results are presented in Fig. 6. When $\alpha=1$, the model showed similar performance to training without SKD, as the distillation loss $L_{\text{skd}}$ was not applied. As $\alpha$ decreased below 1, incorporating $L_{\text{skd}}$ led to improved accuracy, confirming the benefit of SKD. Regarding $\beta$, the best performance was observed around $\beta=0.9$. However, setting $\beta$ above 0.9 resulted in slight performance degradation, likely due to over-smoothing effects, which hinder the model’s responsiveness to new gradients. For the temperature parameter $\tau$, the effect on the classification accuracy was minimal across a wide range of values, indicating that the model is relatively insensitive to $\tau$ when sufficiently trained.
@@ 4.5 Analysis of Self-knowledge Distillation Methods
To evaluate the effectiveness and flexibility of our framework, we conducted experiments that combine our dual-model weight selection method with multiple SKD techniques. Using ViT-S as the teacher model and ViT-T as the student model, experiments were conducted across chest X-ray, lung CT, and brain MRI datasets. Table 6 summarizes the classification accuracy for each method.
We compared the following three SKD strategies: (1) basic SKD, (2) class similarity-based knowledge distillation (CS-KD) [54], and (3) teacher-free knowledge distillation (TF-KD) [33]. CS-KD enhances traditional SKD by modifying the soft labels using a class similarity matrix, allowing the student model to learn from both the correct and closely related semantic classes. This is especially useful in medical image tasks where the class boundaries are subtle. In contrast, TF-KD uses the student model's outputs from previous epochs as pseudo teachers. By referencing past predictions, the model encourages consistency and smooth knowledge transfer across training. Both outputs were softened using temperature scaling, and their KL divergence was used as a loss term.
Overall, all methods combined with weight selection achieved consistently high classification accuracy, confirming the robustness of our approach. Our results demonstrate that the proposed method remains effective regardless of the specific SKD strategy applied. This suggests that the proposed framework is highly adaptable and can be adapted to various distillation strategies. The consistent improvements over CM1 also highlight that combining weight selection with SKD effectively enhances the learning ability of small models, especially in low-data medical imaging scenarios.
@@ 4.6 Analysis of Weight Selection Methods
4.7 Computational Cost and Training Time Analysis
Table 8: Experimental results of the computational cost and training time. 1% of the training data was used.
Dataset Method Cost (MiB) Time (s)
X-ray Ours 2,434 1,328
CM1 1,932 1,031
CT Ours 2,620 1,584
CM1 2,210 1,255
MRI Ours 2,620 916
CM1 2,378 670
We evaluated the computational cost and training time of the proposed method compared to CM1 across three datasets: chest X-ray, lung CT, and brain MRI. To maintain consistency, all experiments used ViT-S as the teacher model and ViT-T as the student model. Table 8 summarizes the GPU memory consumption and training duration for both methods.
Although the proposed method incurs a slight increase in GPU and training time compared to CM1, this overhead is minimal and manageable. In return, it consistently achieves a substantially higher classification accuracy across all datasets. This trade-off highlights the efficiency of our approach by integrating dual-model weight selection with SKD, which substantially boosts model performance without imposing excessive computational cost.
@@ 5 Discussion
A key strength of our method is its dual-model architecture, where two lightweight student models are initialized with distinct, complementary subsets of weights extracted from a large pretrained teacher model. This design enables broad and diverse knowledge transfer while maintaining computational and memory overhead. Moreover, by incorporating SKD, the main model learns not only from ground truth labels but also from the soft predictions of the auxiliary model, enhancing feature learning and generalization. Our results also confirm the framework's adaptability, showing that various SKD strategies—such as CS-KD and TF-KD—can be seamlessly integrated to boost performance. Additionally, swapping the roles of the main and auxiliary models causes negligible variation in results, highlighting the method’s stability and architectural robustness. Overall, the proposed method offers a scalable and practical solution for high-accuracy, resource-efficient medical image classification. Its strong performance under constrained data and hardware conditions makes it a promising candidate for deployment in clinical and point-of-care settings.
Nonetheless, certain limitations remain. Our evaluations are confined to three types of imaging modalities, and generalization to other medical domains—such as ultrasound, endoscopy, or histopathology—needs exploration. Further, a detailed study on how the capacity and architecture of the teacher model influence student model performance could provide valuable insights for optimizing transfer strategies.

### v2

@@ 1 Introduction
Deep learning has become indispensable in the medical field, driving significant advances in diagnostic imaging, disease prediction, and electronic health record analysis [1, 2]. Large-scale models have achieved remarkable success across diverse tasks; however, their deployment in clinical settings is often hindered by limited computational resources [3, 4]. For example, analyzing high-resolution medical images such as chest X-rays and magnetic resonance imaging (MRI) scans requires substantial computing power and memory, making large model deployment impractical in many healthcare environments [5, 6, 7]. Simultaneously, the growing shortage of healthcare professionals is increasing the need for efficient automated systems to reduce workloads [8, 9]. Consequently, there is a rising demand for compact models that can operate effectively under resource constraints while maintaining performance comparable to large-scale counterparts.
A critical factor in improving small-model performance is weight initialization. Poor initialization can lead to vanishing or exploding gradients, hindering effective training. While commonly used methods such as Xavier [10] and Kaiming [11] initialization support stable training from scratch, pretrained large-scale models (e.g., ImageNet-21K [12] or LAION-5B [13]) are increasingly adopted for transfer learning. However, fine-tuning these large models is often infeasible in clinical settings because it demands high-end GPUs with large memory and substantial computing resources—conditions rarely met in real-world medical environments. To address this, recent studies have explored weight selection methods, which transfer only a subset of weights from a large pretrained model to a smaller one [14, 15, 16].
@@ 2.2 Knowledge Distillation
SKD eliminates the need for an external teacher by enabling the model to improve itself. Zhang et al. [45] introduced a method for segmenting deep models such as ResNet into blocks, allowing intra-model knowledge transfer between components. Kim et al. [46] introduced progressive SKD, and Ge et al. [47] developed batch-wise ensembling of self-predictions to suppress incorrect predictions. Recent work has adapted these ideas for transformers [48] and even explored self-distillation in the absence of ground-truth labels using self-supervised techniques [49]. MiniLM [50] exemplifies a task-agnostic SKD method that maintains performance while significantly reducing the model size. Beyond image classification, KD and SKD have also been extended to more complex visual tasks. In 3D vision, Cao et al. [51] formulated unpaired shape completion as a knowledge-transfer problem using a teacher–assistant–student framework. In remote sensing, Du et al. [52] injected land-use knowledge into a multiscale CNN for single-image height estimation, while Li et al. [53] proposed a patch-similarity-based self-distillation scheme that improves cross-view geo-localization by distilling multi-scale local cues. These applications demonstrate that distillation can effectively enhance representation learning under domain shifts and structural ambiguities, which is closely related to our goal of robust medical image classification. A comprehensive survey on knowledge distillation summarizes these trends and outlines the future directions [54].
@@ 3 Dual-model Weight Selection and Self-knowledge Distillation
Our method employs a pretrained large model as the teacher and small-scale models as students, integrating dual-model weight selection for effective knowledge transfer and SKD for efficient training. First, weight selection transfers essential parameters from the teacher model to the student models, enabling them to inherit prior knowledge and achieve strong early performance. Then, SKD allows the student models to refine their learning by leveraging their own predictions, improving task performance, and maximizing the benefit of the initial weights without imposing a high computational cost. Traditional pruning or saliency-driven subset selection aims primarily at parameter reduction, often at the cost of representational diversity. In contrast, our method derives two complementary parameter subsets that strictly preserve structural index alignment, ensuring that both networks retain coherent functional mappings while encouraging diversified feature specialization. Moreover, rather than relying on a separate and typically larger teacher model as in standard teacher–student distillation, we adopt an EMA-updated auxiliary student as a self-adaptive teacher. This design provides stable supervision signals and prevents training collapse, while avoiding additional backpropagation overhead, thus achieving both efficiency and performance gains.
@@ 3.2 Self-knowledge Distillation
where $\eta$ denotes the learning rate, and $\theta_{S}$ represents the parameters of the main student model $S$. Then, the weights of the auxiliary student model $S^{\prime}$ are updated using an exponential moving average (EMA), in which the parameters of model $S$ are integrated into $S^{\prime}$, with the stop gradient (SG) operation applied. EMA and SG are applied during training to stabilize the learning process and reduce update variance. Although $S^{\prime}$ is updated from $S$ via EMA, the two models retain complementary parameter subsets, which leads them to develop distinct yet structurally aligned feature representations. This differentiation reduces the likelihood of mutual error amplification. Moreover, the EMA update introduces temporal averaging, which stabilizes the supervisory signal by preventing abrupt changes in $S^{\prime}$, thereby functioning as a regularization mechanism rather than direct weight replication. After each training iteration, the weights of the model $S^{\prime}$ are updated using:
@@
Table 5: Experimental results of image classification on three full datasets.
Model Dataset Ours CM1
ViT X-ray 0.990 0.949
CT 0.994 0.992
MRI 0.992 0.989
ConvNeXt X-ray 0.956 0.942
CT 0.995 0.994
MRI 0.998 0.996
Table 6: Experimental results of image classification on three extremely limited (0.5%) training datasets.
Model Dataset Ours CM1
ViT X-ray 0.679 $\pm$ 0.012 0.672 $\pm$ 0.015
CT 0.695 $\pm$ 0.010 0.692 $\pm$ 0.011
MRI 0.610 $\pm$ 0.027 0.602 $\pm$ 0.028
Table 7: Evaluation using other metrics with a 10% training data of each dataset.
Model Dataset Ours CM1
ACC F1 Score AUC ACC F1 Score AUC
ViT X-ray 0.866 $\pm$ 0.003 0.878 $\pm$ 0.003 0.968 $\pm$ 0.001 0.839 $\pm$ 0.006 0.842 $\pm$ 0.008 0.958 $\pm$ 0.002
CT 0.897 $\pm$ 0.008 0.884 $\pm$ 0.010 0.962 $\pm$ 0.003 0.843 $\pm$ 0.010 0.846 $\pm$ 0.011 0.943 $\pm$ 0.003
MRI 0.844 $\pm$ 0.012 0.837 $\pm$ 0.014 0.964 $\pm$ 0.003 0.828 $\pm$ 0.007 0.818 $\pm$ 0.008 0.959 $\pm$ 0.004
ConvNeXt X-ray 0.909 $\pm$ 0.003 0.920 $\pm$ 0.002 0.982 $\pm$ 0.001 0.838 $\pm$ 0.011 0.839 $\pm$ 0.014 0.962 $\pm$ 0.017
CT 0.925 $\pm$ 0.004 0.931 $\pm$ 0.004 0.981 $\pm$ 0.003 0.906 $\pm$ 0.002 0.917 $\pm$ 0.002 0.972 $\pm$ 0.001
MRI 0.900 $\pm$ 0.015 0.895 $\pm$ 0.015 0.980 $\pm$ 0.005 0.835 $\pm$ 0.010 0.827 $\pm$ 0.010 0.958 $\pm$ 0.004
@@ 4.1 Dataset and Settings
4.2 Main Results
Tables 2, 3, and 4 present the classification results for chest X-ray, lung CT scan, and brain MRI images, respectively. Each experiment was repeated five times, and both the mean and standard deviation were recorded. For each dataset, training subsets corresponding to 1%, 5%, and 10% of the full dataset were constructed. To minimize sampling bias and maintain comparability, stratified random sampling was employed to preserve the original class distribution. The sampling procedure was repeated five times with distinct random seeds, yielding five independent experimental splits for each subset ratio. All models were trained separately on these splits, and the mean and standard deviation of the performance metrics are reported, ensuring a robust evaluation of the method’s stability under different data availability conditions. These tables compare the performance of CM1, CM2, CM3, and CM4 with the proposed method, which combines weight selection and SKD (Ours).
Our method consistently outperformed all comparison methods in classification accuracy across all datasets, as shown in Tables 2, 3, and 4. By integrating dual-model weight selection with SKD, it enables efficient knowledge transfer and reuse, significantly improving feature retention and learning efficiency. Compared to CM1, which applies weight selection alone, our approach significantly improves feature retention and learning efficiency. Moreover, by evaluating different initialization strategies, we observed that our method substantially enhances neural network initialization, yielding more effective training than CM2, CM3, and CM4. The consistent superiority across datasets demonstrates the robustness and generalizability of our approach, even in low-data regimes. Notably, the significant improvements over CM2–CM4 emphasize that default or classical initialization schemes alone are insufficient for effective learning in smaller models. Our method not only boosts accuracy but also ensures stability across trials, as reflected in the low standard deviations reported. Figures 3, 4, and 5 illustrate the confusion matrices of our method on the chest X-ray, lung CT scan, and brain MRI datasets, respectively. The results demonstrate consistently high accuracy and robust generalization across all modalities. In the chest X-ray dataset (Fig. 3), the model accurately distinguishes COVID-19 from other classes, between clinically similar conditions such as viral pneumonia and lung opacity, underscoring its robustness even with limited data. In the lung CT dataset (Fig. 4), the model shows accurate classification, especially between COVID-19 and normal cases, which are often difficult to separate due to radiological differences. For the brain MRI dataset (Fig. 5), the model effectively identifies tumor types such as glioma and meningioma, while reducing misclassification involving the no tumor class. These results confirm that combining dual-model weight selection with SKD enhances feature representation and discrimination, even when trained on just 10% of the available data.
Furthermore, additional experiments were conducted using the full dataset to provide a more complete comparison with conventional fine-tuning. As shown in Table 5, results from the full dataset confirmed the effectiveness of this method even with large-scale data. In particular, our method achieved higher classification accuracy in the chest X-ray dataset, which includes a larger number of images and more classes, compared to other datasets, demonstrating a strong generalization capability for complex data. On the other hand, since the performance improvement was more pronounced in the limited-data setting, the proposed method is especially effective when training data are scarce.
Additionally, Table 6 summarizes the results obtained using only 0.5% of the training data. These findings further highlight the advantage of our method in extremely data-limited scenarios. Even with such a minimal amount of training samples, the proposed approach consistently outperformed conventional fine-tuning methods, demonstrating strong robustness against data scarcity. This result reinforces that the proposed framework is particularly suitable for medical imaging applications, where annotated data are often minimal. To ensure a more thorough evaluation, we supplemented accuracy with the F1-score and AUC, particularly given the class imbalance present in medical imaging datasets. These metrics were computed under the same experimental protocol applied in Figures 35. As shown in Table 7, our method achieves superior performance across all three metrics, demonstrating consistent advantages in both classification effectiveness and decision boundary quality.
4.3 Analysis of Interchanged Main and Auxiliary Models
4.4 Analysis of Hyperparameters
We conducted experiments using ViT-S as the teacher model and ViT-T as the student model to investigate the influence of three key hyperparameters: $\alpha$, $\beta$, and $\tau$. The results are summarized in Fig. 6. When $\alpha=1$, the model exhibited performance comparable to training without SKD, since the distillation loss $L_{\text{skd}}$ was effectively excluded. As $\alpha$ decreased below 1, the inclusion of $L_{\text{skd}}$ consistently improved accuracy, demonstrating the effectiveness of self-knowledge distillation. For the EMA momentum $\beta$, the highest performance was achieved around $\beta=0.9$, whereas larger values slightly degraded accuracy due to over-smoothing, which can limit the model’s adaptability to new gradient updates. Regarding the temperature parameter $\tau$, classification accuracy remained relatively stable across a broad range of values, suggesting that the model is insensitive to $\tau$ once sufficient training convergence is reached.
@@ 4.5 Analysis of Self-knowledge Distillation Methods
Table 11: GPU memory cost, training time per epoch, computational complexity (FLOPs), and inference speed (FPS) comparison. The training data used was a 1% X-ray dataset.
Method Memory (M) Time (s) FLOPs (G) FPS (fps)
Ours 2,434 3.24 1.078 201.44
CM1 1,932 2.69 1.078 167.10
To evaluate the effectiveness and flexibility of our framework, we conducted experiments that combine our dual-model weight selection method with multiple SKD techniques. Using ViT-S as the teacher model and ViT-T as the student model, experiments were conducted across chest X-ray, lung CT, and brain MRI datasets. Table 9 summarizes the classification accuracy for each method. We compared the following three SKD strategies: (1) basic SKD, (2) class similarity-based knowledge distillation (CS-KD) [61], and (3) teacher-free knowledge distillation (TF-KD) [37]. CS-KD enhances traditional SKD by modifying the soft labels using a class similarity matrix, allowing the student model to learn from both the correct and closely related semantic classes. This is especially useful in medical image tasks where the class boundaries are subtle. In contrast, TF-KD uses the student model's outputs from previous epochs as pseudo teachers. By referencing past predictions, the model encourages consistency and smooth knowledge transfer across training. Both outputs were softened using temperature scaling, and their KL divergence was used as a loss term. Overall, all methods combined with weight selection achieved consistently high classification accuracy, confirming the robustness of our approach. Our results demonstrate that the proposed method remains effective regardless of the specific SKD strategy applied. This suggests that the proposed framework is highly adaptable and can be adapted to various distillation strategies. The consistent improvements over CM1 also highlight that combining weight selection with SKD effectively enhances the learning ability of small models, especially in low-data medical imaging scenarios.
@@ 4.6 Analysis of Weight Selection Methods
Figure 7: Grad-CAM visualization results on CT images. From left to right, three correctly classified samples and one misclassified sample are shown. For each image, the upper label indicates the classification result, and the lower label indicates the ground truth. Red regions correspond to areas where the model exhibits close attention. In successful cases, the model accurately focused on lesion-related or anatomically relevant regions, whereas in the misclassified case, attention was diffusely distributed across irrelevant areas, indicating that the model failed to capture discriminative features.
4.7 Computational Efficiency Analysis
We compared the computational cost and training efficiency of the proposed method and CM1 on the chest X-ray dataset. To ensure a fair comparison, both methods employed ViT-S as the teacher model and ViT-T as the student model. Table 11 presents the GPU memory consumption, training time per epoch, theoretical computational complexity (FLOPs), and inference throughput (FPS).
The proposed method introduces a modest increase in GPU memory usage and a slightly longer training time per epoch. However, this overhead is minor and remains well within a practical range. Also, the method consistently achieves higher classification accuracy on the chest X-ray dataset. In terms of inference efficiency, both methods share identical theoretical FLOPs, indicating that they require the same nominal computation. The slight reduction in FPS observed in our method is due to the auxiliary-model interaction introduced by the self-knowledge distillation process, which incurs minimal runtime overhead without altering the model architecture. Nonetheless, the inference speed remains comparable to CM1, demonstrating that the proposed method maintains real-world deployability while providing improved predictive performance.
4.8 Visualization of Misclassified Samples
To further examine model behavior beyond aggregate performance metrics, we analyzed several misclassified cases using Grad-CAM, as shown in Fig. 7. CT images were selected for visualization due to their clearer structural contrast, which facilitates the interpretation of the highlighted regions. In successful classifications, the model consistently focused on lesion-related or anatomically relevant regions, reflecting meaningful feature localization. For example, in CAP cases, attention was concentrated in bright lesion regions, while for Normal and COVID-19 cases, the model emphasized darker lung areas. In contrast, misclassified samples exhibited diffuse or misplaced attention, indicating difficulty in identifying discriminative structural cues. These observations suggest that failure cases may arise from low-contrast regions or overlapping texture patterns, where the distinction between classes becomes less pronounced, and the model’s decision boundaries are less reliable. Notably, we did not observe consistent associations between these misclassifications and specific weight subset configurations, suggesting that the errors arise primarily from image-level ambiguity rather than the dual-model selection or distillation strategy itself.
@@ 5 Discussion
Nonetheless, certain limitations remain. First, while our method demonstrates robustness across three imaging modalities, its performance under significant domain shifts such as unseen scanner types and varying acquisition protocols has not been fully explored. Future work will examine the performance of the proposed method under domain shift conditions, such as unseen scanner types and variations in acquisition protocols, to further assess its robustness in real-world clinical environments. Next, in real-world clinical datasets, annotation noise arising from inter-observer variability, ambiguous findings, and report-derived labels is almost unavoidable and can degrade model performance. Although we do not explicitly simulate label noise in our experiments, two components of the proposed framework are expected to confer a certain degree of robustness: (i) the dual-model weight selection favors parameter configurations that generalize better on held-out data, implicitly reducing overfitting to mislabeled samples; and (ii) the SKD stage leverages soft self-predictions as complementary supervision, which can smooth out isolated noisy labels and encourage consistency across related classes. A more systematic study under controlled noise levels and realistic noise patterns remains an important direction for future work. Also, we report the mean ± standard deviation over five independently sampled training subsets to assess performance stability and robustness. While these results show consistent improvements over the baseline, we acknowledge that formal statistical significance testing could further strengthen the conclusions and plan to address this in future work. Additionally, we plan to extend our comparative analysis to include more advanced lightweight architectures, such as MobileNet, EfficientNet, and distilled transformer variants, to more clearly position the proposed method within the broader landscape of contemporary lightweight deep learning approaches. Furthermore, we acknowledge that excessive correlation between the main and auxiliary student models may reduce the benefit of complementary representation learning, and exploring additional diversification strategies constitutes an important direction for future work.